ishita invested a sum of Rs 12000 at 5 percent per annum interest. She received an amount of Rs 13230 after n years. find the value of n.

Answer by Theo(13342)

formula is f = p * (1+r)^n

f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

in your problem:

f = 13230
p = 12000
r = .05
n = what you want to find.

the formula becomes 13230 = 12000 * (1.05)^n

divide both sides of the equation by 12000 to get:
13230/12000 = 1.05^n
take the log of both sides of the equation to get:
log(13230/12000) = log(1.05^n)
since log(1.05^n) is equal to n * log(1.05), your equation becomes:
log(13230/12000) = n*log(1.05)
divide both sides of the equation by log(1.05) to get:
log(13230/12000) / log(1.05) = n
solve for n to get:
n = 2

the money will grow from 12000 to 13230 in 2 years at 5% interest per year.

12000 * 1.05^2 = 13230